a U-Tube manometer has a mercury of relative density of 13.6 as it monumetric liquid. kerosene of relative density of 0.6 was poured into one of the arms and so a drop of 2.0 cm result in that arm. Also, water with a density of 1 g per cm cube was poured into another arm and so the two arms become the same again. calculate the height of the kerosene
1
Expert's answer
2021-07-19T09:51:32-0400

"\text {Let A be the point of the upper level of mercury}"

"\text{shoulders in the pipe where kerosene is injected}"

"\text{Let B be the point of the upper level of mercury}\newline
\text{ of the other arm of the tube}"

"P_A = P_B"

"P_A =\rho_kgh_k"

"\text{where }"

"\rho_k- \text{density of kerosene}"

"h_k-\text{kerosene column height}"

"P_B =\rho_mg\Delta h"

"\text{where}"

"\rho_m - \text {mercury density }"

"\Delta h- \text{difference in mercury levels at points A and B}"

"\Delta h = 2cm \text{ by the condition of the problem}"


"P_A = P_B"

"\rho_kgh_k = \rho_mg\Delta h"

"h_k =\frac{ \rho_m\Delta h}{\rho_k}=\frac{ 13.6*2}{0.6}=45.3cm"


"\text{Answer: }h_k =45.3cm"